If a plane travels at 240km/h for 20 seconds, how far does it go?

Answer by jim_thompson5910(35256)

20 seconds * (1 min/60 seconds)*(1 hr/60 min) = 20*(1/60)*(1/60) = 1/180 hour

20 seconds = 1/180 hour

-------------------------------------------------------

D = r*t

D = 240*(1/180)

D = 4/3

D = 1.33333

The distance it travels is approximately 1.33333 km